Find the general solution of y"-y'-2y=4x2

1
Expert's answer
2021-05-07T11:20:01-0400

Solution

To get general solution of homogeneous equation y0"-y0'-2y0=0 let’s solve characteristic equation r2-r-2=0.

r1,2 = (1±√(1+4*2))/2 = (1±3)/2 => r1 = -1, r2 = 2

Therefore y0(x) = C1e-x+C2e2x    

Partial solution y1(x) may be find in the form of polynomial y1(x) = Ax2+Bx+C. Substituting into equation we will get

2A-2Ax-B-2Ax2-2Bx-2C = 4x2  =>  

2A-B-2C = 0; -2A-2B = 0; -2A = 4  => 

A = -2; B = 2; C = -3

So y1(x) = -2x2+2x-3

General solution of the given equation is

y(x) = y0(x)+ y1(x) =  C1e-x+C2e2x -2x2+2x-3

Answer

y(x) = C1e-x+C2e2x -2x2+2x-3